### v1

# Exact Chiral Symmetries of 3+1D Hamiltonian Lattice Fermions

Abstract
We construct Hamiltonian models on a 3+1d cubic lattice for a single Weyl fermion and for a single Weyl doublet protected by exact (as opposed to emergent) chiral symmetries. In the former, we find a not-on-site, non-compact chiral symmetry which can be viewed as a Hamiltonian analog of the Ginsparg-Wilson symmetry in Euclidean lattice models of Weyl fermions. In the latter, we combine an on-site $U(1)$ symmetry with a not-on-site $U(1)$ symmetry, which together generate the $SU(2)$ flavor symmetry of the doublet at low energies, while in the UV they generate an algebra known in integrability as the Onsager algebra. This latter model is in fact the celebrated magnetic Weyl semimetal which is known to have a chiral anomaly from the action of $U(1)$ and crystalline translation, that gives rise to an anomalous Hall response - however reinterpreted in our language, it has two exact $U(1)$ symmetries that gives rise to the global $SU(2)$ anomaly which protects the gaplessness even when crystalline translations are broken. We also construct an exact symmetry-protected single Dirac cone in 2+1d with the $U(1)\rtimes T$ parity anomaly. Our constructions evade both old and recently-proven no-go theorems by using not-on-siteness in a crucial way, showing our results are sharp.
I Introduction
Regulating chiral gauge theories like the standard model on the lattice has been a long standing problem. As well as offering a route for extracting more precise predictions from these theories, this has been highlighted as a deep theoretical problem thanks to early no-go results [1, 2]. These prove under certain assumptions that these lattice theories always come with fermions of both handedness, a phenomenon known as fermion doubling. This problem already exists for fermions with chiral global symmetries, and can be understood as a ’t Hooft anomaly [3]. More precisely, a system with an on-site global symmetry, meaning one which does not mix degrees of freedom at separate sites, must be free of ’t Hooft anomalies. This rules out, for example, a lattice model with an on-site $U(1)$ global symmetry giving rise to a single charged Weyl fermion in the infrared (IR).
There have been several approaches so far attempted to circumvent the fermion doubling problem. A simple strategy is to try to gap out the unwanted fermions. Because of the above, any such perturbation also breaks the chiral symmetry. However, it was observed by Ginsparg and Wilson [4] that in Euclidean spacetime lattice models, there is a kind of remnant of the chiral symmetry. This is in fact another symmetry [5], which takes the form
$\begin{gathered}\delta\psi=\gamma^{5}(1-\frac{1}{2}aD)\psi,\end{gathered}$ (1)
where $a$ is the lattice spacing, and $D$ is a finite difference operator which is a discretization of the spacetime Dirac operator. This is an example of a not-on-site symmetry, since $(D\psi)(x)$ is a (linear) function of $\psi$ at nearby spacetime lattice points, and not of $\psi(x)$ alone. When $a\to 0$ this becomes the continuum axial symmetry of a Dirac fermion. It thus protects the fermion from perturbatively gaining a mass.
In this paper, we will construct Hamiltonian models on a 3d cubic lattice with chiral global symmetry. These models are Weyl semimetals [6, 7] protected by not-on-site symmetries, and have the advantage over Euclidean lattice models of being realizable in quantum simulators [8, 9].
In particular, we will discuss a model which in the IR gives rise to a single Weyl fermion, and in the ultraviolet (UV) has a not-on-site symmetry which flows to the chiral $U(1)$ symmetry, and protects the Weyl fermion from obtaining a Majorana mass.
Then we will discuss a model containing two Weyl fermions, which is in fact a simple model of a magnetic Weyl semimetal, and show that it has not-on-site symmetries which flow to the chiral $SU(2)$ symmetry of the infrared. This protects both Weyls from obtaining a mass, thanks to Witten’s $SU(2)$ anomaly [10]. This insight offers a very different interpretation of the usual chiral anomaly involving $U(1)$ and translation [11, 12, 13], as such an anomaly holds even when crystalline translation symmetry is broken. We develop a formalism which allows one to explore general not-on-site symmetries acting linearly on the fermion operators, analogous to the Ginsparg-Wilson symmetry.
Finally, we will discuss time reversal symmetry in these systems, and build a 2+1d model with a single massless Dirac fermion in the IR, protected by the parity anomaly.
II Symmetry-protected single Weyl fermion in 3+1d
Here we build a (time-reversal broken) 3+1d tight-binding model with finite-range hopping and a single Weyl node at crystalline momentum $\mathbf{k}=\mathbf{0}$ that is protected by a finite-range non-on-site chiral symmetry. We start with a two-band model (two fermion species per site, which we think of as spin $s\in\{\uparrow,\downarrow\}$) on a cubic lattice, known as a magnetic Weyl semimetal, described by the second-quantized Hamiltonian
$H_{2}=\sum_{\mathbf{k},s,s^{\prime}}c^{\dagger}_{\mathbf{k},s}h_{2}(\mathbf{k})_{ss^{\prime}}c_{\mathbf{k},s^{\prime}}\quad,$ (2)
where $h_{2}(\mathbf{k})$, the Bloch Hamiltonian, is given by
$h_{2}(\mathbf{k})=$ $\sin k_{x}\sigma^{x}+\sin k_{y}\sigma^{z}+\left[\sin k_{z}+m(\mathbf{k})\right]\sigma^{y}\quad,$ (3)
with $m(\mathbf{k})=2-\cos k_{x}-\cos k_{y}$, and $\boldsymbol{\sigma}=(\sigma^{x},\sigma^{y},\sigma^{z})$ are Pauli matrices acting on the spin degree of freedom. It hosts two Weyl fermions at low energy, seen by linearizing the Hamiltonian are the node at momentum $\mathbf{k}_{1}\equiv\mathbf{0}$ and another at momentum $\mathbf{k}_{2}\equiv(0,0,\pi)$.
This model has an on-site $U(1)$ symmetry $c_{\mathbf{r}}\mapsto e^{-i\theta\hat{Q}_{0}}c_{\mathbf{r}}e^{i\theta\hat{Q}_{0}}=e^{i\theta}c_{\mathbf{r}}$, generated by the on-site charge $\hat{Q_{0}}=\sum_{\mathbf{r},s}c_{\mathbf{r},s}^{\dagger}c_{\mathbf{k},s}$. We will need to break this in order to gap the Weyl node at $\mathbf{k}_{2}$. In order to do this, let us write our Hamiltonian in the BdG formalism with the basis $d^{\dagger}_{\mathbf{k}}\equiv(c^{\dagger}_{\mathbf{k}\uparrow},c^{\dagger}_{\mathbf{k}\downarrow},c_{-\mathbf{k}\uparrow},c_{-\mathbf{k}\downarrow})$, such that the Hamiltonian takes the form
$h_{2}^{\mathrm{BdG}}(\mathbf{k})=$ $\frac{1}{2}\left[\sin k_{x}\sigma^{x}+\sin k_{y}\sigma^{z}+\sin k_{z}\tau^{z}\sigma^{y}+m(\mathbf{k})\sigma^{y}\right],$ (4)
[…]
$\tau^{x}h^{\mathrm{BdG}}(\mathbf{k})^{T}\tau^{x}=-h^{\mathrm{BdG}}(-\mathbf{k})\quad.$ (5)
In this formalism the $U(1)$ symmetry is not automatic, it is generated by $\tau^{z}$. We can add a $U(1)$ breaking term such as $(1-\cos k_{z})\tau^{x}\sigma^{y}$ to gap the Weyl node at $\mathbf{k}_{2}$, leading to the modified Hamiltonian
$h_{\text{single Weyl}}^{\mathrm{BdG}}(\mathbf{k})=$ $\frac{1}{2}\,[\sin k_{x}\sigma^{x}+\sin k_{y}\sigma^{z}+m(\mathbf{k})\sigma^{y}$
$+\sin k_{z}\tau^{z}\sigma^{y}+(1-\cos k_{z})\tau^{x}\sigma^{y}]\,.$ (6)
This Hamiltonian has a single Weyl node remaining at $\textbf{k}_{1}=0$, and no other gapless modes.
It turns out $h_{\text{single Weyl}}^{\mathrm{BdG}}(\mathbf{k})$ commutes with a symmetry generator $S_{\text{chiral}}(\mathbf{k})$ given by
$S_{\text{chiral}}(\mathbf{k})=\frac{1}{2}\left[(1+\cos k_{z})\tau^{z}+\sin k_{z}\tau^{x}\right]\quad.$ (7)
At $\mathbf{k}_{1}$ the symmetry reduces to $S_{\text{chiral}}(\mathbf{k}_{1})=\tau^{z}$ which is just the original $U(1)$ charge operator for the corresponding single-particle modes. Therefore, it gives an exact chiral symmetry in this model.
We can see by inspection that this symmetry prevents all mass terms, but allows for terms such as $\sigma^{y}$ and $(1+\cos k_{z})\tau^{z}+\sin k_{z}\tau^{x}$ which only shifts the Weyl node in momentum space. Thus, the anomaly indeed stabilizes the low energy theory. One interesting caveat is that at $\mathbf{k}_{2}$, $S_{\text{chiral}}(\mathbf{k}_{2})=0$, which is what allowed us to gap the second Weyl node there. We could eventually symmetrically move the remaining Weyl node to this point as well and then completely gap the system.
We can write the associated charge operator
$\hat{Q}_{\text{chiral}}=\sum_{\mathbf{k}}d_{\mathbf{k}}^{\dagger}\,S_{\text{chiral}}(\mathbf{k})\,d_{\mathbf{k}}\quad,$ (8)
via a Fourier transform in real space as
$\hat{Q}_{\text{chiral}}=\frac{1}{2}\sum_{\mathbf{\mathbf{r}},s}\bigg{[}c_{\mathbf{r},s}^{\dagger}c_{\mathbf{r},s}+c_{{\mathbf{r}+\hat{z}},s}^{\dagger}c_{\mathbf{r},s}-ic_{{\mathbf{r}+\hat{z}},s}^{\dagger}c_{\mathbf{r},s}^{\dagger}\bigg{]}+h.c.$ (9)
We see that, like the Ginsparg-Wilson symmetry, this charge operator involves nearest neighbor couplings, and is thus not-on-site
Moreover, we see that the BdG generator obeys $S_{\text{chiral}}(\mathbf{k})^{2}=\cos^{2}(k_{z}/2)\mathds{1}$. Thus, it has a continuous spectrum, and is therefore a non-compact symmetry, generating an $\mathbb{R}$ action on the full Hilbert space. This is necessary to evade the no-go theorem in [14], which proved that a locality-preserving chiral $U(1)$ symmetry of a single Weyl fermion does not exist. Curiously, as a Hamiltonian itself, it describes decoupled wires along the $z$-axis of massless Majorana fermions. [^1] A non-Hermitian symmetry generator of a single Weyl fermion was proposed in [15], by interpolating between the identity and a translation symmetry.
We can give an alternative proof of this no-go result, as follows, which applies more generally to the types of symmetries we have been considering. The idea is to observe that $\hat{Q}_{\text{chiral}}$ itself may be viewed as a Hamiltonian with chiral symmetry. If $\hat{Q}_{\text{chiral}}$ could be chosen to have a quantized spectrum, then since it is a two band model, after suitable normalization, one band would be at chiral charge $+1$ and the other at $-1$. As a Hamiltonian then, $\hat{Q}_{\text{chiral}}$ would describe a half-filled band insulator with anomalous chiral $U(1)$ symmetry, which is impossible.
III Symmetry-protected double Weyl fermion in 3+1d
The symmetry (7), which has continuous spectrum, may be naturally separated into two generators with quantized spectrum:
$\begin{gathered}\hat{S}_{0}(\textbf{k})=\tau^{z}\qquad\hat{Q}_{0}=\sum_{\mathbf{k}}d_{\mathbf{k}}^{\dagger}\hat{S}_{0}(\mathbf{k})d_{\mathbf{k}}\\ \hat{S}_{1}(\textbf{k})=\cos k_{z}\tau^{z}+\sin k_{z}\tau^{x}\qquad\hat{Q}_{1}=\sum_{\mathbf{k}}d_{\mathbf{k}}^{\dagger}\hat{S}_{1}(\mathbf{k})d_{\mathbf{k}}.\end{gathered}$ (10)
The first is the usual $U(1)$ symmetry, while the second is composed of Kitaev Majorana chains [16] along $z$-axis wires. These generators do not commute, instead they generate an infinite-dimensional Lie algebra known as the Onsager algebra, introduced in [17]. This algebra has recently appeared in the study of the 1+1D chiral anomaly on the lattice [18, 19], and offers another route to defining exact symmetries on the lattice giving anomalous symmetries in the IR.
We can actually write a Hamiltonian that has this symmetry and two Weyl nodes, which in the BdG formalism above (using the $d_{\mathbf{k}}^{\dagger}$ basis) is
$h_{\text{double Weyl}}^{\mathrm{BdG}}(\mathbf{k})=$ $\mathds{1}_{\tau}\otimes\frac{1}{2}\bigg{[}\sin k_{x}\sigma^{x}+\sin k_{y}\sigma^{z}$
$+\left[\cos k_{z}-\cos K+m(\mathbf{k})\right]\sigma^{y}\bigg{]}\quad,$ (11)
where the identity in the $\tau$ basis ensures it has both $U(1)$ symmetries, $K$ is a parameter, and $m(\textbf{k})$ is the same as in (3). This model is a magnetic Weyl semimetal model that has the two Weyl nodes at $\textbf{k}=\pm\mathbf{K}$ where $\mathbf{K}=(0,0,K)$.
Let’s linearize around the Weyl nodes. We get
$\begin{gathered}h^{\text{BdG}}_{l}=\mathds{1}_{\tau}\otimes\frac{1}{2}(k_{x}\sigma^{x}+k_{y}\sigma^{z}-\sin K\ k_{z}\sigma^{y}).\end{gathered}$ (12)
This shows that for $K\neq 0,\pi$, the two Weyl nodes have an opposite handedness. To figure out the effect of $\hat{S}_{1}(\textbf{k})$ at low energy, we can also linearize it, and obtain
$\begin{gathered}\hat{S}_{1,K}(\textbf{k})=\cos K\ \tau^{z}+\sin K\ \tau^{x}.\end{gathered}$ (13)
The important feature for $K\neq 0,\pi$ is that the second term is non-zero.
[…]
We can demonstrate that this symmetry protects the gapless Weyl points. To do so, we must break translation symmetry, since otherwise $z$-axis translations also act as a discrete axial symmetry and help to stabilize the Weyl nodes [12, 13]. To analyze translation-symmetry breaking, we consider an extended basis $e_{\mathbf{k}}\equiv(c_{\mathbf{k}-\mathbf{K}},c^{\dagger}_{-\mathbf{k}+\mathbf{K}},c_{\mathbf{k}+\mathbf{K}},c^{\dagger}_{-\mathbf{k}-\mathbf{K}})^{T}$ (we suppress the spin component). Hamiltonians in this basis may couple states at $\mathbf{k}$ with $\mathbf{k}+2\mathbf{K}$ but are automatically $\hat{Q}_{0}$ preserving. In this basis the symmetry action of $\hat{Q}_{1}$ becomes
$U(1)_{\hat{Q}_{1}}:\delta e_{\mathbf{k}}=$ $i(\cos k_{z}\cos K\ \tau^{z}+\sin k_{z}\sin K\ \,\eta^{z}\tau^{z}$
$+$ $\sin k_{z}\cos K\ \tau^{x}-\cos k_{z}\sin K\ \,\eta^{z}\tau^{x})e_{\mathbf{k}},$ (14)
which prohibits all mass terms except $m_{j}(\mathbf{k})\eta^{z}\sigma^{j}$. However, these terms always commute with at least one term in the original Hamiltonian, so the result is a shift in the gapless modes rather than a gap. At a large enough pertubation, we can move the modes until $K=0$ or $\pi$, where the symmetry generators are aligned and no longer generate the whole $SU(2)$ symmetry. At these special points, we will be able to open a symmetric gap.
IV Time Reversal Symmetric Single Weyl Fermion in 3+1d
So far, we have considered time-reversal breaking models. We can also construct time-reversal invariant models, at the cost of making the symmetry generator slightly more not-on-site. As long as $S(\textbf{k})$ is a smooth function of the momentum, then the charge density in real space will be a sum of terms with faster-than-polynomial decay. Such “almost-local” operators share many properties with local operators, while being closed under Hamiltonian evolution generated by such terms [20, 21]. This will allow us to employ bump functions and partitions of unity in momentum space.
To construct a time-reversal invariant model with a single protected Weyl fermion, we begin with a model on a cubic lattice with eight Weyl nodes. We use the BdG formalism with the basis $d^{\dagger}_{\mathbf{k}}\equiv(c^{\dagger}_{\mathbf{k}\uparrow},c^{\dagger}_{\mathbf{k}\downarrow},c_{-\mathbf{k}\uparrow},c_{-\mathbf{k}\downarrow})$ used above, giving the Hamiltonian
$h^{\mathrm{BdG}}_{8}(\mathbf{k})=\frac{1}{2}\left[\sin k_{x}\sigma^{x}+\sin k_{y}\sigma^{z}+\sin k_{z}\tau^{z}\sigma^{y}\right].$ (15)
This model has Weyl nodes at all eight time-reversal-invariant-momentum (TRIM) points of the Brillouin zone, as well as a time-reversal symmetry $\Theta=i\sigma^{y}\mathcal{K}$, where $\mathcal{K}$ is complex conjugation, satisfying $\Theta^{2}=-1$.
We will now add a $U(1)$ symmetry-breaking term that will gap out all Weyl nodes except the one at $\mathbf{k=0}$. In order to facilitate our discussion, let us first define a bump function $B(k)$ given by
$B(\mathbf{k},w)=\begin{cases}e^{-\frac{w^{2}}{|\mathbf{k}|^{2}-w^{2}}}&\mathrm{for}\,\,|\mathbf{k}|<w\\ 0&\mathrm{for}\,\,|\mathbf{k}|\geq w\end{cases}$ (16)
where $w>0$ determines the width of the bump. This function is smooth but non-analytic. We add a $U(1)$ breaking term such that the total Hamiltonian is now given by
$h^{\mathrm{BdG}}_{\text{TRS Weyl}}(\mathbf{k})=$
$h_{8}^{\text{BdG}}(\textbf{k})+\sum_{j\in\{x,y,z\}}$ $B\left(k_{j}-\pi,\frac{\pi}{2}\right)(1-\cos k_{j})\tau^{y}\sigma^{y}\quad,$ (17)
which gaps all Weyl nodes except the one at $\mathbf{k}=0$. By inspection, this preserves the time-reversal symmetry $\Theta$. It also has an almost-local chiral symmetry generator
$S_{\text{chiral}}(\mathbf{k})=B\left(\mathbf{k},\frac{\pi}{2}\right)\tau^{z}\quad.$ (18)
As previously, this chiral symmetry is not quantized, as it must be by [14] and our arguments in the previous section. We could also choose a step function instead of a bump function for this symmetry, and get a quantized chiral symmetry, but in real space it would not be almost-local, with the charge density having algebraic decay. This also avoids [14] because such an operator does not generate a locality preserving unitary evolution.
V Parity anomaly of a Single Dirac fermion with time-reversal symmetry in 2+1d
We start with a 2+1d time-reversal invariant Dirac fermion model on a square lattice with four Dirac nodes given by the BdG Hamiltonian with basis $d^{\dagger}_{\mathbf{k}}\equiv(c^{\dagger}_{\mathbf{k}\uparrow},c^{\dagger}_{\mathbf{k}\downarrow},c_{-\mathbf{k}\uparrow},c_{-\mathbf{k}\downarrow})$
$h_{4}^{\text{BdG}}(\mathbf{k})=\mathds{1}_{\tau}\otimes\frac{1}{2}(\sin k_{x}\sigma^{x}+\sin k_{y}\sigma^{y})\quad.$ (19)
This model has Dirac nodes are at all four TRIM points of the Brillouin zone, and a time-reversal symmetry $\Theta=i\sigma^{y}\mathcal{K}$ with $\Theta^{2}=-1$. We will now add a $U(1)$ symmetry-breaking term that will gap out all Dirac nodes except the one at $\mathbf{k=0}$:
$h^{\mathrm{BdG}}_{\text{single Dirac}}(\mathbf{k})=$ (20)
$h_{4}^{\text{BdG}}(\textbf{k})+$ $\sum_{j\in x,y}B\left(k_{j}-\pi,\frac{\pi}{2}\right)(1-\cos k_{j})\tau^{y}\sigma^{y},$
This gaps all Dirac nodes except the one at $\mathbf{k}=0$. This model commutes with an almost-local symmetry generator of the same form as (18):
$S(\mathbf{k})=B\left(\mathbf{k},\frac{\pi}{2}\right)\tau^{z},$ (21)
which commutes with time-reversal $\Theta$ and again has continuous spectrum. Since at $\mathbf{k}=0$, it acts as $\tau^{z}$, it gives rise to the $U(1)$ symmetry of the single Dirac fermion in the IR, which together with $\Theta$ protects this Dirac fermion from gaining a mass by the parity anomaly. Note that, analogous to the previous examples, by a large enough perturbation, we can push the Dirac cone into the region where $S(\textbf{k})=0$ and eventually gap out the system.
As with the 3+1d chiral symmetry, this particle number symmetry of this type—acting linearly on fermions and commuting with a time-reversal action—must be non-quantized. This follows the same argument as in 3+1d. Otherwise, we could consider the $U(1)$ symmetry generator itself as a $U(1)\rtimes T$ symmetric Hamiltonian. For these two band models, it would describe a band insulator with a parity anomaly, which is impossible.
VI Discussion
In this work, we have introduced several Hamiltonian models with new, not-on-site symmetries giving rise to anomalous symmetries acting on the low energy fermions. We have realized the chiral anomaly of a single charged Weyl fermion in 3+1d, the $SU(2)$ anomaly of a doublet of left-handed Weyl fermions in 3+1d, and the $U(1)\rtimes T$ parity anomaly of a single Dirac fermion in 2+1d.
The chiral symmetry in 3+1d has a nearest-neighbor charge operator, but is non-quantized, as it must be by the no-go theorem of [14] and our arguments above. The two $SU(2)$ generators we constructed are quantized (and one is on-site), but don’t satisfy the expected $SU(2)$ Lie algebra, instead satisfying an Onsager algebra. This is consistent with the anomaly since either $U(1)$ symmetry on its own is anomaly-free. In this way, it is structurally similar to the $U(1)_{V}\times U(1)_{A}$ anomaly realized in 1+1D in [18, 19], where both generators are quantized and one is on-site, but they don’t commute.
To include time reversal symmetry in these systems, we had to relax our charge density operators to be almost-local operators, with faster-than-polynomial decaying tails. We are unsure if this is a necessary further weaking of on-siteness, but it is convenient.
In each example, we are able to show a no-go theorem that shows our construction is nearly as good as possible. The method for proving these no-go theorems seems very general for studying anomalous $U(1)$ symmetries. If we have a $U(1)$ symmetry generator $\hat{Q}$, which is either anomalous or shares an anomaly with other symmetries commuting with it. Then we can regard $\hat{Q}$ itself as a Hamiltonian with these anomalous symmetries, including $\hat{Q}$. Therefore, $\hat{Q}$ must have a non-trivial ground state.
There is a fun example which avoids this no-go argument in 2+1d. Let us take a 2d square lattice of spin-$\frac{1}{2}$ degrees of freedom, and take $\hat{Q}$ to be the toric code Hamiltonian [22], which has an integer spectrum. It also commutes with time reversal given by complex conjugation, and together, these two symmetries generate a bosonic $U(1)\rtimes T$ parity anomaly [23]. The ground state of the toric code Hamiltonian is indeed non-trivial, and sufficient to match this parity anomaly.
Since Weyl semimetals are physically realizable in a variety of condensed matter experiments [24, 25] it may be interesting to explore potentially observable effects of this low-energy $SU(2)$ symmetry and associated anomaly in such systems, although the global nature of the anomaly may complicate things.
Acknowledgements: LG acknowledges support from the Walter Burke Institute for Theoretical Physics at Caltech and the Caltech Institute for Quantum Information and Matter. RT acknowledges support from the Mani L. Bhaumik Presidential Term Chair in Theoretical Physics at UCLA.